OpenStudy (anonymous):

The sum of the first 6 terms of the geometric sequence 1, -2, 4, -8, ... is?

OpenStudy (anonymous):

multiply your sequence by -2 each time. (1*-2) =-2 -2*-2=4 4*-2= -8 -8*-2= 16 16*-2= -32 now you can do the sum